OpenStudy (bohotness):

What is the solution to y = 2x – 1 for the replacement set x = {1.4}? ( , ) can you please explan?

OpenStudy (bohotness):

@TheAsker2002

OpenStudy (bohotness):

@iGreen

OpenStudy (bohotness):

@paki

OpenStudy (igreen):

Plug in 1.4 for 'x'

OpenStudy (igreen):

y = 2x - 1 y = 2(1.4) - 1

OpenStudy (igreen):

2 * 1.4 = ?

OpenStudy (bohotness):

2.8

OpenStudy (igreen):

Yes, and 2.8 - 1 = ?

OpenStudy (bohotness):

1.8

OpenStudy (igreen):

Yes, so we have y = 1.8, so our answer is (1.4, 1.8).

OpenStudy (bohotness):

those are my answers?

OpenStudy (igreen):

Points are in the form of (x, y)..and we were given x, 1.4 in the question, and we got 'y' when we put in 1.4 for 'x' in the equation.

OpenStudy (igreen):

Yes, they are.

OpenStudy (bohotness):

:D

OpenStudy (bohotness):

thank you

OpenStudy (igreen):

Happy to help